OpenStudy (anonymous):

Two trains each make a journey of 330km. One of the train travels 5km/h faster than the other and takes 30 min less time. Find the speed of the trains

OpenStudy (dan815):

read that wwrong lol

OpenStudy (whpalmer4):

train 1 has speed \(s_1\). train 2 has speed \(s_2 = s_1+5\). train 1 takes time \(t_1\) to travel 330 km. train 2 takes time \(t_2 = t_1-\frac{1}{2}\) to travel 330 km. \(d = st\) is the formula for distance traveled in time \(t\) at speed \(s\). We can write two equations for the trains: \[330 \text{ km} = s_1*t_1\]\[330 \text{ km} = s_2*t_2 = (s_1+5)*(t_1-\frac{1}{2})\]

OpenStudy (whpalmer4):

Solve those two for \(s_1,t_1\) and then use the relationship between \(s_1,s_2\) to get the other train's speed.

OpenStudy (dan815):

t*(x+5)=330 = x*(t+0.5) note the 30 mins became 0.5 hours

OpenStudy (whpalmer4):

right, that's an essential point!

OpenStudy (whpalmer4):

You get two solutions, but only one of them will make sense.